# Roughening dynamics of interfaces in two-dimensional quantum matter

### v1

@@ Introduction.
The static and dynamical properties of interfaces are fundamental to understand the physics and to engineer the functionality of materials. Interfaces can even undergo their own unique phase transitions independent of the bulk matter [1, 2, 3, 4, 5, 6, 7]. At these roughening transitions, the nature of interfaces changes qualitatively from being smooth to rough [8]. This is not only relevant for many-body physics but also in a broader scientific context such as for instance in high-energy physics, where the flux tube connecting quarks realizes an interface that can undergo roughening [9]. For quantum matter, roughening transitions can also occur at zero temperature in 2D systems with discrete symmetry such as in the quantum Ising model [10]. The dynamical exploration of such interface roughening in 2D has remained an outstanding challenge to date.
In this work we study the non-equilibrium signatures of the roughening transition in quantum matter. For that purpose we perform large-scale Tree Tensor Network simulations for the real-time evolution of an initial domain wall state in a 2D quantum Ising model on a square lattice [11, 12, 13]. We find that the underlying quantum roughening transition reflects in a qualitative change of dynamical behavior of the domain wall state upon tuning the transverse-field strength. Below the roughening transition for weak transverse fields we observe long-lived prethermal plateaus, which we quantitatively explain with an effective quantum solid-on-solid (SOS) model. Above the roughening transition, but still within the ferromagnetic phase, we find that the domain wall profile decays quickly, consistent with a roughening of the domain wall interface. The dynamical scenario investigated here can be readily explored experimentally in Rydberg atomic systems [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
@@
where the first sum runs over all neighboring spin pairs. The TFIM exhibits a ferromagnetic phase, that extends to non-zero temperatures, with a quantum phase transition at $g_{c}/J\approx 3.04$ and a thermal transition temperature of $T_{c}/J\approx 2.27$ at $g=0$ [29]. Importantly, the 2D quantum Ising model inherits a second quantum phase transition point of Berezinski-Kosterlitz-Thouless (BKT) type within the ferromagnetic phase, which is associated with a transition from smooth to rough interfaces [10]. In the quantum domain, when considering the limit $J\gg g$, the TFIM can be perturbatively linked to the PXP model, which shows strong dynamical constrains and Hilbert space fragmentation, leading to a slow relaxation of various domain wall initial conditions [30, 31, 13, 12].
In the following we depart from any such perturbative limit and target the dynamics in strongly correlated regimes. For the purpose of studying the dynamics of quantum interfaces, we will initialize the system with two oppositely polarized magnetic domains, separated by a straight domain wall, see Fig. 1a. As will be shown later, this class of initial conditions leads to prethermal plateaus in the regime of smooth interfaces, which cannot be explained by Hilbert space fragmentation. Instead, this non-perturbative effect is related to the thermalization of the interface below a roughening temperature.
@@
Figure 1: Time evolution of a flat interface on an 8$\times$8 lattice with open boundary conditions. (a) Spatially resolved magnetizations at times $tJ=0$, $10$ and $100$ for a transverse field of $g/J=0.75$. The bottom three plots show (b) the imbalance, (c) domain wall length, and (d) entanglement entropy for several transverse fields. The imbalance shows the existence of long-lived plateaus, even at transverse fields $g\approx J$. All the results are shown for three different bond dimensions $\chi=181,256,362$, where the opacity increases with the bond dimension. For panel (a) and (b), almost all of the data points lie on top of each other.
@@
Fig. 1 includes a convergence check of the numerical results with different bond dimensions up to $\chi=362$. While late times and stronger transverse fields become challenging, the prominent pre-thermal plateaus are well within the regime of certain convergence. Most sensitive to varying bond dimension is the entanglement entropy (Fig. 1d), which shows that the moderate amounts of entanglement generated up to the lifetime of the plateaus are well captured.
@@
Figure 2: (a) Mapping from a domain wall state in the full two dimensional model to an effective one-dimensional bosonic state. (b) Expectation value of the kink operator, evaluated on the groundstate of the effective model as a function of the transverse field. Different maximal occupation numbers $4\leq N_{\mathrm{max}}\leq 14$ are shown; as $N_{\mathrm{max}}$ increases, the kink operator starts exhibiting a jump at $g_{R}/J\approx 1.38$ corresponding to the roughening transition, see also the SM [38]. (c) Expectation value of the kink operator, evaluated on the thermal state of the effective model in the classical limit ($g=0$) as a function of temperature. We consider chain lengths $L_{x}\in[500,512\cdot 10^{3}]$, with exponentially increasing steps. The inset shows that the crossover temperature $T_{R}$, defined by $\langle K(T_{R}/J)\rangle=0.5$, vanishes in the thermodynamic limit as the inverse logarithm of the system size, confirming the analytical prediction (5) which is indicated by the dashed grey line. The constant $b$ in the x-axis label is given by $b=2(1-\cos 1)/\log 2$. Data points in (b) and (c) are represented by crosses, with a linear interpolation between them. (d) Phase diagram of the 2D TFIM. The blue line represents the critical line separating the ferro- from the paramagnetic phase, the red line denotes the effective temperature of the initial state based on its energy, obtained from QMC simulations, see SM [38]. Our results point towards the existence of a roughening transition at a value of $g_{R}/J\approx 1.38<g_{c}/J$. The green shaded region shows the extended smooth interface regime, present in finite systems.
In order to explain the observed non-perturbative prethermal effects, we formulate an effective, one-dimensional model which describes domain wall states without overhangs or bubbles in the bulk,
@@
Here, we introduced height operators $N_{i}$ measuring the perpendicular displacement of the domain wall, see Fig. 2a, and raising (lowering) operators $E_{i}^{\dagger}$ ($E_{i}$). The raising and lowering operators, obey the commutation relations $\left[E_{i},N_{j}\right]=E_{i}\delta_{i,j}$ similar to the standard definition used in quantum rotor models [39]. Hence, the phase operator $\varphi$ defined through $E_{j}=\exp(i\varphi_{j})$ is canonically conjugate to the height operator and the $N_{i}$ can alternatively be viewed as bosonic occupation number operators. It is important to note that the model is not based on some low-order Schrieffer-Wolff transformation, but instead is formulated such that it captures the relevant (vertical) fluctuations of the initial domain wall.
In solid-on-solid-like (SOS) models as defined in Eq. (3), roughening appears as a Berezinskii–Kosterlitz–Thouless (BKT) transition [40]. The critical point of $\mathrm{H}_{\mathrm{eff}}$ has been argued to be upper-bounded by the critical point of the quantum rotor model [41, 42, 43]. This is noticeably below the symmetry-breaking phase transition of the full TFIM.
@@
The effective model covers the subspace of single domain-wall states without bubbles or overhang. To check the validity of this description, we plot the time-dependence of the horizontal domain wall length $D_{x}=1/2\sum_{i,j}(\operatorname{\mathds{1}}-S^{x}_{i,j}S^{x}_{i,j+1})$ in Fig. 3a. The small deviations of $\langle D_{x}\rangle/L_{x}$ from one for transverse fields up to $g/J\approx 1$ support the validity of the effective model in that regime.
@@
Figure 3: (a) Time evolution of the horizontal contribution to the domain wall length divided by the horizontal lattice dimension. The plot shows that up until $g/J=1$, each column has approximately one horizontal domain wall for the times considered, confirming the validity of the effective model (3) in that regime. (b) Comparison between the effective (dashed lines/crosses) and the full model (solid lines/circles) for the time evolution of the kink operator. In order to filter out the strong fluctuations of the data we show the running mean of the kink operator. For the full model, we calculate the modified kink operator $K_{\mathrm{M}}=\langle K\rangle/\langle K_{\mathrm{bulk}}\rangle$. We don’t show data for transverse fields beyond $g/J=1$, as bulk and interface contributions become increasingly difficult to disentangle. (c) Time averages of the kink operators shown in (b) as functions of the transverse field. The values agree well up to $g/J=1$, once again confirming the quantitative predictive power of the effective model in that regime.
We find very good quantitative agreement between these two models, meaning that the dynamics of interfaces in the TFIM can be understood in terms of the SOS model: For the considered intermediate values of the transverse field, the SOS model thermalizes within the smooth interface regime (cf. Fig. 2d) and the initially flat domain wall profile remains eternally stable. This stability of the domain wall manifests itself in the full TFIM in the form of prethermal plateaus with $\bar{K}_{\mathrm{M}}(t)>0$. In contrast to the effective model, a subsequent decay of these prethermal states is to be expected in the full model, but these timescales seem out of reach for current numerical approaches. Fig. 3c displays the late-time stationary values of the kink operator, which once more highlight the compelling quantitative agreement between the SOS model and the TFIM. Finally, notice that the imbalance in Fig. 1b and the kink operator in Fig. 3b exhibit different lifetimes of the prethermal plateaus. This indicates that restoring the rotational symmetries requires longer times than the restoration of translational symmetry.
@@ Acknowledgements.
We acknowledge fruitful discussions with J.-D. Urbina, M. Steinhuber, L. Behringer, S. Tomsovic, T. Mendes-Santos, A. Gambassi, U.-J. Wiese.
The tree tensor network simulations presented in this work were produced with TTN.jl [54], a software package we developed based on the ITensor library [55]. The MPS simulations were performed with the help of ITensor. The QMC simulations were produced with the help of the ALPS software package [56]. MS and WK were supported through the Helmholtz Initiative and Networking Fund, Grant No. VH-NG-1711. We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769 (NT, MR and MS), under Grant No. 277101999 – CRC network TR 183 (NT and MR), under project 499180199 – FOR 5522 (MH). MH has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 853443). MR and NT acknowledge funding under Horizon Europe programme HORIZON-CL4-2022-QUANTUM-02-SGA via the project 101113690 (PASQuanS2.1). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS [57] and through FZJ on JURECA [58] at Jülich Supercomputing Centre (JSC). Data and code will be made available soon.
@@ Appendix A End Matter
$\braket{K_{\alpha}(L_{x})}\approx\exp\left(-2(1-\cos(\alpha)L_{x}\right)\,.$ (9)
@@
Supplemental materials for “Roughening dynamics of interfaces in two-dimensional quantum matter”
@@ Appendix B Tree Tensor Networks and Time Dependent Variational Principle
Another consequence of the binary tree structure is the fact that the computational cost of tensor contractions is of $\mathcal{O}(\chi^{4})$, whereas the QR decomposition costs $\mathcal{O}(\chi^{3})$. This means, that the overall runtime of serial algorithms is dominated by tensor contractions and, therefore, the simulations profit substantially from GPU acceleration. In our case, we observe a speedup of around 70×70$\times$70 × over CPU performance. More specifically, the most expensive simulation at a bond dimension $\chi=362$, at a time step $dt=0.1$ up to $t_{\mathrm{max}}=100$ takes around two days on a single A100 GPU – on a CPU (Xeon Gold) the same calculation would take around 145 days.
Appendix C Kink operator

### v2

# Roughening dynamics of interfaces in the two-dimensional quantum Ising model

@@ Introduction.
The static and dynamical properties of interfaces are fundamental to understand the physics and to engineer the functionality of materials. Interfaces can even undergo their own unique phase transitions independent of the bulk matter. At a roughening transition, first identified in models for surface growth and classical magnetism [1, 2, 3], the nature of interfaces changes qualitatively from being smooth to rough. Intuitively these different phases can be characterized by their interface fluctuations, with small and bounded fluctuations in the smooth phase and large, and unbound ones in the rough phase. The theoretical analysis of effective surface-on-surface (SOS) models revealed the Berezinski-Kosterlitz-Thouless (BKT) nature of the transition [4, 5], which was confirmed in numerical simulations [6]. For experimental studies, the interface between liquid and crystallized Helium-4 is a particularly suited model system [7]. At the involved low temperatures, however, quantum effects enter. It was found that quantum fluctuations can cause a roughening transition at vanishing temperature $T=0$ in two-dimensional systems, but not in three dimensions [8, 9]. Interface properties are relevant also in a broader scientific context beyond condensed matter. For instance, in high-energy physics, the flux tube connecting quarks realizes an interface that can undergo roughening [10]. Signatures of which were recently observed via digital quantum simulation of a lattice gauge theory [11]. The exploration of such interface roughening in quantum matter away from equilibrium has, however, remained an outstanding challenge to date.
In this work we study the non-equilibrium signatures of the roughening transition in the quantum Ising model. Based on large-scale Tree Tensor Network simulations for real-time evolution [12, 13, 14], we find that the underlying quantum roughening transition is reflected in a qualitative change of dynamical behavior of domain wall initial conditions upon tuning the transverse field strength. In particular, we identify the independent equilibration of the interface in the smooth interface regime as an alternative cause of prethermal plateaus, which is distinct from the known mechanisms relying on approximate conservation laws or Hilbert space fragmentation [15, 16]. The phenomenology can be readily explored in experiments with Rydberg atomic systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
where the first sum runs over all neighboring spin pairs. The TFIM exhibits a ferromagnetic phase, that extends to non-zero temperatures, with a quantum phase transition at $g_{c}/J\approx 3.04$ and a thermal transition temperature of $T_{c}/J\approx 2.27$ at $g=0$ [32]. Importantly, the 2D quantum Ising model inherits a second quantum phase transition point of BKT type within the ferromagnetic phase, which is associated with a transition from smooth to rough interfaces [9]. A suited order parameter will be defined at a later point of this work, providing a quantitative description of the transition. In the quantum domain, when considering the limit $J\gg g$, the TFIM can be perturbatively linked to the PXP model, which shows strong dynamical constrains and Hilbert space fragmentation, leading to a slow relaxation of various domain wall initial conditions [33, 34, 14, 13].
For the purpose of studying the dynamics of quantum interfaces, we will initialize the system with two oppositely polarized magnetic domains, separated by a straight domain wall, see Fig. 1a. This is a natural choice for experimental platforms, while not introducing any additional effects stemming from e.g. interface curvature. For weak enough transverse fields, i.e., within the PXP approximation, these states remain stable due to sectors of different interface lengths being dynamically disconnected from each other, preventing thermalization up to timescales which scale exponentially with $J/g$  [^1] Eventual thermalization is expected to happen for transverse fields $g/J\geq 0.2$ as is revealed by an analysis of the levelspacing statistics of a $4\times 5$ system, see SM for more details [35]..
In the following we depart from any such perturbative limit and target the dynamics in strongly correlated regimes $g/J\sim 1$. As will be shown later, we observe pre-thermal plateaus which cannot be explained by Hilbert space fragmentation. Instead, this non-perturbative effect is related to the thermalization of the interface below a roughening temperature.
@@
Figure 1: Time evolution of a flat interface on an 8$\times$8 lattice with open boundary conditions. (a) Spatially resolved magnetizations at times $tJ=0$, $10$ and $100$ for a transverse field of $g/J=0.75$. The bottom three plots show (b) the imbalance, (c) the domain wall length, and (d) the entropy of entanglement across the initial interface for several transverse fields: the color coding is indicated in the color bar to the right. The imbalance shows the existence of long-lived plateaus, even at transverse fields $g\approx J$. All the results are shown for three different bond dimensions $\chi=181,256,362$, where the opacity increases with the bond dimension. For panel (a) and (b), almost all of the data points lie on top of each other.
@@
Fig. 1 includes a convergence check of the numerical results with different bond dimensions up to $\chi=362$. While late times and stronger transverse fields become challenging, the prominent pre-thermal plateaus are well within the regime of certain convergence. Most sensitive to varying bond dimension is the half-system entanglement entropy (Fig. 1d), when splitting the system into two equal partitions along the initial interface, which shows that the moderate amounts of entanglement generated up to the lifetime of the plateaus are well captured.
@@
Figure 2: (a) Mapping from a domain wall state in the full two dimensional model to an effective one-dimensional bosonic state. (b) Expectation value of the kink operator, evaluated on the groundstate of the effective model as a function of the transverse field. Different maximal occupation numbers $4\leq N_{\mathrm{max}}\leq 14$ in steps of two are shown; as $N_{\mathrm{max}}$ increases, the kink operator starts exhibiting a jump at $g_{R}/J\approx 1.38$ corresponding to the roughening transition, see also the SM [35]. (c) Expectation value of the kink operator, evaluated on the thermal state of the effective model in the classical limit ($g=0$) as a function of temperature. We consider chain lengths $L_{x}\in[500,512\cdot 10^{3}]$, with exponentially increasing steps. The inset shows that the crossover temperature $T_{R}$, defined by $\langle K(T_{R}/J)\rangle=0.5$, vanishes in the thermodynamic limit as the inverse logarithm of the system size, confirming the analytical prediction (5) which is indicated by the dashed grey line. The constant $b$ in the x-axis label is given by $b=2(1-\cos 1)/\log 2$. Data points in (b) and (c) are represented by crosses, with a linear interpolation between them. (d) Phase diagram of the 2D TFIM. The blue line represents the critical line separating the ferro- from the paramagnetic phase, the red line denotes the effective temperature of the initial state based on its energy, obtained from QMC simulations, see SM [35]. Our results point towards the existence of a roughening transition at a value of $g_{R}/J\approx 1.38<g_{c}/J$. The green shaded region shows the extended smooth interface regime, present in finite systems.
In order to explain the observed non-perturbative prethermal effects, we formulate an effective, one-dimensional model which corresponds to the projection of Eq. (1) onto the subspace of domain wall states with exactly one horizontal interface segment per column (see SM for a more detailed derivation [35]),
@@
Here, we introduced height operators $N_{i}$ measuring the perpendicular displacement of the domain wall, see Fig. 2a. The raising (lowering) operators $E_{i}^{\dagger}$ ($E_{i}$) are the projection of the $\sigma_{i}^{z}$, which in the manifold of single horizontal domain walls can only flip spins next to the domain wall. The raising and lowering operators, obey the commutation relations $\left[E_{i},N_{j}\right]=E_{i}\delta_{i,j}$ similar to the standard definition used in quantum rotor models [45]. Hence, the phase operator $\varphi$ defined through $E_{j}=\exp(i\varphi_{j})$ is canonically conjugate to the height operator and the $N_{i}$ can alternatively be viewed as bosonic occupation number operators. It is important to note that the model is not based on some low-order Schrieffer-Wolff transformation, but instead captures the relevant fluctuations of the initial domain wall.
The effective model (3) is also closely related to solid-on-solid-like (SOS) models [4, 5], in which roughening appears as a Berezinskii–Kosterlitz–Thouless (BKT) transition [46]. The critical point of $\mathrm{H}_{\mathrm{eff}}$ has been argued to be upper-bounded by the critical point of the quantum rotor model [47, 48, 49]. This is noticeably below the symmetry-breaking phase transition of the full TFIM.
@@
The effective model covers the subspace of single domain-wall states without bubbles or overhang. To check the validity of this description, we plot the time-dependence of the horizontal domain wall length $D_{x}=1/2\sum_{i,j}(\operatorname{\mathds{1}}-\sigma^{x}_{i,j}\sigma^{x}_{i,j+1})$ in Fig. 3a. The small deviations of $\langle D_{x}\rangle/L_{x}$ from one for transverse fields up to $g/J\approx 1$ support the validity of the effective model in that regime.
@@
Figure 3: (a) Time evolution of the horizontal contribution to the domain wall length divided by the horizontal lattice dimension. The plot shows that up until $g/J=1$, each column has approximately one horizontal domain wall for the times considered, confirming the validity of the effective model (3) in that regime. (b) Comparison between the effective (dashed lines/crosses) and the full model (solid lines/circles) for the time evolution of the kink operator. In order to filter out the strong fluctuations of the data we show the running mean of the kink operator. For the full model, we calculate the modified kink operator $K_{\mathrm{M}}=\langle K\rangle/\langle K_{\mathrm{bulk}}\rangle$. We don’t show data for transverse fields beyond $g/J=1$, as bulk and interface contributions become increasingly difficult to disentangle. (c) Late-time averages of the kink operators shown in (b), taken over the interval $tJ\in[20,100]$, as functions of the transverse field. The values agree well up to $g/J=1$, once again confirming the quantitative predictive power of the effective model in that regime.
We find very good quantitative agreement between these two models, meaning that the dynamics of interfaces in the TFIM can be understood in terms of the SOS model: For the considered intermediate values of the transverse field, the SOS model thermalizes within the smooth interface regime (cf. Fig. 2d) and the initially flat domain wall profile remains eternally stable. This stability of the domain wall manifests itself in the full TFIM in the form of prethermal plateaus with $\bar{K}_{\mathrm{M}}(t)>0$. In contrast to the effective model, a subsequent decay of these prethermal states is to be expected in the full model, but these timescales seem out of reach for current numerical approaches. Fig. 3c displays the late-time stationary values of the running, time-averaged kink operator by taking the mean over points lying in the interval $tJ\in[20,100]$, which once more highlight the compelling quantitative agreement between the SOS model and the TFIM. Results obtained for a 16 by 16 lattice confirm said agreement for larger system sizes, without the need to suppress temporal fluctuations via taking the running mean, see SM [35] for more details. Finally, notice that the imbalance in Fig. 1b and the kink operator in Fig. 3b exhibit different lifetimes of the prethermal plateaus. This indicates that restoring the rotational symmetries requires longer times than the restoration of translational symmetry.
@@ Acknowledgements.
We acknowledge fruitful discussions with A. Segner, J.-D. Urbina, M. Steinhuber, L. Behringer, S. Tomsovic, T. Mendes-Santos, A. Gambassi, U.-J. Wiese. The tree tensor network simulations presented in this work were produced with TTN.jl [58], a software package we developed based on the ITensor library [59]. The MPS simulations were performed with the help of ITensor. The QMC simulations were produced with the help of the ALPS software package [60]. MS and WK were supported through the Helmholtz Initiative and Networking Fund, Grant No. VH-NG-1711. We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769 (NT, MR and MS), under Grant No. 277101999 – CRC network TR 183 (NT and MR), under project 499180199 – FOR 5522 (MH). MH has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 853443). MR and NT acknowledge funding under Horizon Europe programme HORIZON-CL4-2022-QUANTUM-02-SGA via the project 101113690 (PASQuanS2.1). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS [61] and through FZJ on JURECA [62] at Jülich Supercomputing Centre (JSC). Data and code are available at [63].
@@ Appendix A End Matter
$\braket{K_{\alpha}(L_{x})}\approx\exp\left(-2q(1-\cos(\alpha)L_{x}\right)\,.$ (9)
@@
Supplemental materials for “Roughening dynamics of interfaces in the two-dimensional quantum Ising model”
@@ Appendix B Tree Tensor Networks and Time Dependent Variational Principle
Another consequence of the binary tree structure is that the tensors appearing in the tree are of size $\chi^{2}\times\chi$ and thus of the skinny-tall type, which allows for an efficient parallelization of the QR decompositions on these tensors. Since tensor contractions are already naturally parallelizable, this means that the overall runtime of the simulations profit substantially from GPU acceleration. In our case, we observe a speedup of around 70×70$\times$70 × over CPU performance. More specifically, the most expensive simulation at a bond dimension $\chi=362$, at a time step $dt=0.1$ up to $t_{\mathrm{max}}=100$ takes around two days on a single A100 GPU – on a CPU (Xeon Gold) the same calculation would take around 145 days.
Appendix C Effective model
In the following section we present and motivate a derivation of the effective model. The domain wall length operator, introduced in the main text, can be split into a horizontal and vertical component, by performing the sum only over horizontally
$D_{x}=\frac{1}{2}\sum_{i,j}\left(\operatorname{\mathds{1}}-\sigma^{x}_{i,j}\sigma^{x}_{i,j+1}\right)$ (S1)
or vertically
$D_{y}=\frac{1}{2}\sum_{i,j}\left(\operatorname{\mathds{1}}-\sigma^{x}_{i,j}\sigma^{x}_{i+1,j}\right)$ (S2)
oriented bonds, respectively. With that, the transverse field Ising Hamiltonian can be rewritten as
$\mathrm{H}=2JD_{x}+2JD_{y}-g\cdot\sum_{i}\sigma^{z}_{i}\,,$ (S3)
which coincides with the original Hamiltonian up to an irrelevant, additive constant. The effective model is introduced as the projection onto the subspace of domain wall states that have exactly one horizontal domain wall per column, thus fixing $D_{x}/L_{x}=1$. As a consequence these states cannot exhibit overhangs or bubbles of oppositely oriented spins within the bulk.
The vertical contribution to the domain wall length operator is determined by the difference of the vertical position $N_{i},\,i\in\{1,...,L_{x}\}$ of two neighboring segments
$D_{y}\propto\sum_{i}|N_{i}-N_{i-1}|\,,$ (S4)
while the absolute value reflects that the result is independent of the relative ordering in each pair.
Appendix D Kink operator